\begin{thebibliography}{dlPG99}


%\bibitem{Ab} S. Abdulali, Abelian varieties and the general Hodge conjecture, Comp. Math. 109 (1997), 341---355,

\bibitem{An} Y. Andr\'e, Motifs de dimension finie (d'apr\`es S.-I. Kimura, P. O'Sullivan,...), S\'eminaire Bourbaki 2003/2004, Ast\'erisque 299 Exp. No. 929, viii, 115---145,

%\bibitem{And} Y. Andr\'e, Pour une th\'eorie inconditionnelle des motifs, Publications math\'ematiques de l'IHES 83 no 1 (1996), 5---49,

%\bibitem{A} D. Arapura, Motivation for Hodge cycles, Advances in Math. vol. 207 (2006), 762---781,

%\bibitem{Ay} J. Ayoub, Motives and algebraic cycles: a selection of conjectures and open questions, preprint available from user.math.uzh.ch/ayoub/,

%\bibitem{Beau0} A. Beauville, Some remarks on K\"ahler manifolds with $c_1=0$, in: Classification of algebraic and analytic manifolds (Katata, 1982), Birkh\"auser Boston, Boston 1983,

%\bibitem{Beau} A. Beauville, Sur l'anneau de Chow d'une vari\'et\'e ab\'elienne, Math. Ann. 273 (1986), 647---651,

%\bibitem{Beau2} A. Beauville, Some surfaces with maximal Picard number, Journal de l'Ecole Polytechnique Tome 1 (2014), 101---116,

%\bibitem{Bl2} S. Bloch, Some elementary theorems about algebraic cycles on abelian varieties, Invent. Math. 37 (1976), 215---228,

%\bibitem{B} S. Bloch, Lectures on algebraic cycles, Duke Univ. Press Durham 1980,

%\bibitem{BO} S. Bloch and A. Ogus, Gersten's conjecture and the homology of schemes, Ann. Sci. Ecole Norm. Sup. 4 (1974), 181---202,

%\bibitem{BS} S. Bloch and V. Srinivas, Remarks on correspondences and algebraic cycles, American Journal of Mathematics Vol. 105, No 5 (1983), 1235---1253,

%\bibitem{Bo} C. Borcea, K3 surfaces with involution and mirror pairs of Calabi---Yau manifolds, in: Mirror symmetry,
%II, AMS Studies in Adv. Math. 1, AMS 1997,

\bibitem{Br} M. Brion, Log homogeneous varieties, in: Actas del XVI Coloquio Latinoamericano de Algebra, 
Revista Matem\'atica Iberoamericana, Madrid 2007,
arXiv: math/0609669,

%\bibitem{Bus} N. Buskin, Every rational Hodge isometry between two $K3$ surfaces is algebraic, arXiv:1510.02852v1,

%\bibitem{Cat} F. Catanese, Surfaces with $K^2=p_g=1$ and their period mapping, in:  Algebraic geometry (Copenhagen, 1978), Springer Lecture Notes in Mathematics, Springer 1979,

%\bibitem{CD} Clingher and C. Doran, Note on a geometric isogeny of $K3$ surfaces, Int. Math. Research Notices 2011 (2011), 3657---3687,

%\bibitem{CDP} P. Candelas, X. de la Ossa and L. Parkes, Generalized Calabi--Yau manifolds and the mirror of a rigid manifold, Nucl. Phys. B 406 (1993), 115---154,

%\bibitem{CG} A. Cattaneao and A. Garbagnati, Calabi--Yau $3$--folds of Borcea--Voisin type and elliptic fibrations, arXiv:1312.3481v1,

%\bibitem{CHSW} P. Candelas, G. Horowitz, A. Strominger and E. Witten, Vacuum configurations for superstrings, Nucl. Phys. B 258 (1985), 46---74,

\bibitem{CM} 
M. de Cataldo and L. Migliorini, The Chow groups and the motive of the Hilbert scheme of points on a
surface, Journal of Algebra 251 no. 2 (2002), 824---848,

%\bibitem{ChM} F. Charles and E. Markman, The standard conjectures for holomorphic symplectic varieties deformation equivalent to Hilbert schemes of $K3$ surfaces, Comp. Math. 149 (2013),
%481---494,

\bibitem{Del} P. Deligne, La conjecture de Weil pour les surfaces $K3$, Invent. Math. 15 (1972), 206---226,

%\bibitem{DM} C. Deninger and J. Murre, Motivic decomposition of abelian schemes and the Fourier transform. J. reine u.
%angew. Math. 422 (1991), 201---219,

%\bibitem{Diaz} H. Diaz, The motive of the Fano surface of lines, arXiv:1602.06403v1,

%\bibitem{EL} H. Esnault and M. Levine, Surjectivity of cycle maps, in: Journ\'ees de G\'eom\'etrie alg\'ebrique d'Orsay, Juillet 1992, Ast\'erisque 218 (1993),

%\bibitem{LFu} L. Fu, On the action of symplectic automorphisms on the $CH_0$--groups of some hyper-K\"ahler fourfolds, Math. Z. 280 (2015), 307---334,

%\bibitem{LFu} L. Fu, Decomposition of small diagonals and Chow rings of hypersurfaces and Calabi--Yau complete intersections, Advances in Mathematics (2013), 894---924,

%\bibitem{F} W. Fulton, Intersection theory, Springer--Verlag Ergebnisse der Mathematik, Berlin Heidelberg New York Tokyo 1984,

%\bibitem{GS} S. Galkin and E. Shinder, The Fano variety of lines and rationality problem for a cubic hypersurface, arXiv:1405.5154,

\bibitem{vG} B. van Geemen, Kuga--Satake varieties and the Hodge conjecture, in:
The Arithmetic and Geometry of Algebraic Cycles, Banff 1998 (B. Gordon et alii, eds.), Kluwer Dordrecht 2000,

%\bibitem{vG2} B. van Geemen, Half twists of Hodge structures of CM--type, J. Math. Soc. Japan Vol. 53 No. 4 (2001), 813---833,


\bibitem{GP} V. Guletski\u{\i} and C. Pedrini, The Chow motive of the Godeaux surface, in:
Algebraic Geometry, a volume in memory of Paolo Francia (M.C. Beltrametti et alii, editors),
Walter de Gruyter, Berlin New York, 2002,

%\bibitem{HI} A. Hirschowitz and J. Iyer, Hilbert schemes of fat r--planes and the triviality of Chow groups of complete intersections. In: Vector
%bundles and complex geometry, Contemp. Math. 522, Amer. Math. Soc., Providence 2010,

%\bibitem{Hu} D. Huybrechts, Symplectic automorphisms of $K3$ surfaces of arbitrary order, Math. Res. Letters 19 (2012), 947---951,

\bibitem{Iv} F. Ivorra, Finite dimensional motives and applications (following S.-I. Kimura, P. O'Sullivan and others), in:
        Autour des motifs, Asian-French summer school on algebraic geometry and number theory,
       Volume III, Panoramas et synth\`eses, Soci\'et\'e math\'ematique de France 2011,
          
\bibitem{Iy} J. Iyer, Murre's conjectures and explicit Chow--K\"unneth projectors for varieties with a nef tangent bundle, Transactions of the Amer. Math. Soc. 361 (2008), 1667---1681,

\bibitem{Iy2} J. Iyer, Absolute Chow--K\"unneth decomposition for rational homogeneous bundles and for log homogeneous varieties, Michigan Math. Journal
 Vol.60, 1 (2011), 79---91,
 
%\bibitem{J} U. Jannsen, Mixed motives and algebraic K--theory, Springer Lecture Notes in Mathematics 1400 (1990),

%\bibitem{J} U. Jannsen, 
%Motives, numerical equivalence, and semi-simplicity, Invent. Math. 107(3) (1992), 447---452, 

%\bibitem{J2} U. Jannsen, Motivic sheaves and filtrations on Chow groups, in: Motives (U. Jannsen et alii, editors), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1,  

%\bibitem{J3} U. Jannsen, Equivalence relations on algebraic cycles, in: The arithmetic and geometry
%of algebraic cycles (B. Gordon et alii, eds.), Banff Conference 1998, Kluwer,

\bibitem{J4} U. Jannsen, On finite--dimensional motives and Murre's conjecture, in: Algebraic cycles and motives (J. Nagel and C. Peters, editors), Cambridge University Press, Cambridge 2007,

%\bibitem{KMP} B. Kahn, J. P. Murre and C. Pedrini, On the transcendental part of the motive of a surface, in: Algebraic cycles and motives (J. Nagel and C. Peters, eds.), Cambridge University Press, Cambridge 2007,

\bibitem{KSeb} B. Kahn and R. Sebastian, Smash--nilpotent cycles on abelian 3--folds, Math. Res. Letters 16 (2009), 1007---1010,

\bibitem{KS} T. Katsura and T. Shioda, On Fermat varieties, Tohoku Math. J. Vol. 31 No. 1 (1979), 97---115,

\bibitem{Kim} S. Kimura, Chow groups are finite dimensional, in some sense,
Math. Ann. 331 (2005), 173---201,

%\bibitem{K0} S. Kleiman, Algebraic cycles and the Weil conjectures, in: Dix expos\'es sur la cohomologie des sch\'emas, North--Holland Amsterdam, 1968, 359---386, 

%\bibitem{K} S. Kleiman, The standard conjectures, in: Motives (U. Jannsen et alii, editors), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1, 

%\bibitem{Ko} K. Koike, Elliptic $K3$ surfaces admitting a Shioda--Inose structure, Comment. Math. Univ. St. Pauli 61 No 1 (2012), 77---86,

%\bibitem{Kun} K. K\"unnemann, A Lefschetz decomposition for Chow motives of abelian schemes, Inv. Math. 113 (1993), 85---102,

%\bibitem{moi} R. Laterveer, Algebraic varieties with small Chow groups, J. Math. Kyoto Univ. Vol. 38 No 4 (1998), 673---694,

%\bibitem{yet} R. Laterveer, Yet another version of Mumford's theorem, Archiv Math. (2015),   

%\bibitem{moi2} R. Laterveer, A brief note concerning hard Lefschetz for Chow groups, to appear in Canadian Math. Bulletin, 

%\bibitem{moimult} R. Laterveer, On a multiplicative version of Bloch's conjecture, submitted,

%\bibitem{excubic} R. Laterveer, A family of cubic fourfolds with finite--dimensional motive, preprint,

%\bibitem{exK3} R. Laterveer, A family of $K3$ surfaces with finite--dimensional motive, Arch. Math.,

%\bibitem{blochinv} R. Laterveer, Some results on a conjecture of Voisin for surfaces of geometric genus one, Bol. Unione Mat. Ital.,

%\bibitem{Lew} J. Lewis, Cylinder homomorphisms and Chow groups, Math. Nachr. 160 (1993), 205---221,

%\bibitem{Li} C. Liedtke, Supersingular $K3$ surfaces are unirational, Invent. Math. 200 (2015), 979---1014,

%\bibitem{Ma} S. Ma, On $K3$ surfaces which dominate Kummer surfaces, Proceedings of the American Math. Soc. 141 No 1 (2013), 131---137,

%\bibitem{Mor} D. Morrison, Algebraic cycles on products of surfaces, in: Proceedings, Algebraic Geometry
%Symposium, T\^ohoku University, 1984,

%\bibitem{Mo} D. Morrison, On $K3$ surfaces with large Picard number, Invent. Math. 75 No 1 (1984), 105---121,

%\bibitem{Mor2} D. Morrison, Isogenies between algebraic surfaces with geometric genus one, Tokyo J. Math. Vol 10 no 1 (1987), 179---187,

%\bibitem{Mu} S. Mukai, On the moduli space of bundles on $K3$ surfaces I, in: Vector bundles on algebraic varieties (Bombay, 1984), Tata Institute of Fund. Research, Bombay 1987,

%\bibitem{M} D. Mumford, Rational equivalence of $0$--cycles on surfaces, J. Math. Kyoto Univ. Vol. 9 No 2 (1969), 195---204,

%\bibitem{Mur} J. Murre, On a conjectural filtration on the Chow groups of an algebraic variety, parts I and II, Indag. Math. 4 (1993), 177---201,

\bibitem{MNP} J. Murre, J. Nagel and C. Peters, Lectures on the theory of pure motives, Amer. Math. Soc. University Lecture Series 61, Providence 2013,

%\bibitem{Murt} V. Murty, Algebraic cycles on abelian varieties, Duke Math. J. 50 (1983), 487---504,


%\bibitem{OG} K. O'Grady, Decomposable cycles and Noether--Lefschetz loci, arXiv:1502.07897v2,

%\bibitem{OS} K. Oguiso and J. Sakurai, Calabi---Yau threefolds of quotient type, Asian J. Math. 5 (2001), 43---78,

%\bibitem{Or} D. Orlov, Equivalences of derived categories and $K3$ surfaces, J. Math. Sci. (New York) 84 (1997), 1361---1381,

%\bibitem{OS} P. O'Sullivan, Algebraic cycles on an abelian variety, J. Reine Angew. Math. 654 (2011), 1---81,

%\bibitem{Ot} A. Otwinowska, Remarques sur les groupes de Chow des hypersurfaces de petit degr\'e, C. R.
%Acad. Sci. Paris S\'erie I Math. 329 (1999), no. 1, 51---56,

%\bibitem{Par} K. Paranjape, Abelian varieties associated to certain K3 surfaces, Comp. Math. 68 (1988), 11---22,

\bibitem{P} C. Pedrini, On the finite dimensionality of a $K3$ surface, Manuscripta Mathematica 138 (2012), 59---72,

%\bibitem{Ped} C. Pedrini, Bloch's conjecture and valences of correspondences for $K3$ surfaces, arXiv:1510.05832v1,

%\bibitem{PP} A. del Padrone and C. Pedrini, Derived categories of coherent sheaves and motives of K3 surfaces, in: Regulators (J. Gil et alii, eds.), Contemp. Math. 571, Amer. Math. Soc., Providence RI, 2012,

\bibitem{PW} C. Pedrini and C. Weibel, Some surfaces of general type for which Bloch's conjecture holds, to appear in: Period Domains, Algebraic Cycles, and Arithmetic, Cambridge Univ. Press, 2015,

%\bibitem{PS} C. Peters and J. Steenbrink, Mixed Hodge structures, Springer--Verlag Ergebnisse der Mathematik, Berlin Heidelberg New York 2008,

%\bibitem{RM} J. Ram\'on Mar\'\i, On the Hodge conjecture for products of certain surfaces, Collect. Math. 59, 1 (2008), 1---26,

%\bibitem{Rito} C. Rito, A note on Todorov surfaces, Osaka Journal of Math. 46(3) (2009), 685---693,

%\bibitem{Roh} C. Rohde, Maximal automorphisms of Calabi---Yau manifolds versus maximally unipotent monodromy,
%Manuscripta Math. 131 (2010), 459---474,

%\bibitem{R} A.A. Rojtman, The torsion of the group of 0--cycles modulo rational equivalence, Annals of Mathematics 111 (1980), 553---569,

%\bibitem{Schl} U. Schlickewei, The Hodge conjecture for self--products of certain $K3$ surfaces, Journal of Algebra 324 (2010), 507---529,

%\bibitem{Sch} T. Scholl, Classical motives, in: Motives (U. Jannsen et alii, editors), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1, 

%\bibitem{Seb2} R. Sebastian, Smash nilpotent cycles on varieties dominated by products of curves, Comp. Math. 149 (2013), 1511---1518,

%\bibitem{Seb} R. Sebastian, Examples of smash nilpotent cycles on rationally connected varieties, Journal of Algebra 438 (2015), 119---129,

%\bibitem{SV} M. Shen and C. Vial, The Fourier transform for certain hyperK\"ahler fourfolds, Memoirs of the AMS 240 (2016), no.1139,

\bibitem{Sh} T. Shioda, The Hodge conjecture for Fermat varieties, Math. Ann. 245 (1979), 175---184,

%\bibitem{Tan} S. Tankeev, On the standard conjecture of Lefschetz type for complex projective threefolds. II, Izvestiya Math. 75:5 (2011), 1047---1062,

%\bibitem{T} J. Tate, Algebraic cycles and poles of zeta functions, in: Arithmetical Algebraic Geometry, Harper and Row, New York 1965,

%\bibitem{Tod} A. Todorov, Surfaces of general type with $p_g= 1$ and $(K,K) = 1$, Ann. Sci. de l'Ecole Normale Sup. 13 (1980), 1---21,

%\bibitem{Vial} C. Vial, Pure motives with representable Chow groups, Comptes Rendus de l'Acad\'emie des Sciences 348 (2010), 1191---1195,

%\bibitem{V} C. Vial, Algebraic cycles and fibrations, Documenta Math. 18 (2013), 1521---1553,

\bibitem{V2} C. Vial, Projectors on the intermediate algebraic Jacobians, New York J. Math. 19 (2013), 793---822,

\bibitem{V3} C. Vial, Remarks on motives of abelian type, to appear in Tohoku Math. J.,

%\bibitem{V4} C. Vial, Niveau and coniveau filtrations on cohomology groups and Chow groups, Proceedings of the LMS 106(2) (2013), 410---444,

\bibitem{V5} C. Vial, Chow--K\"unneth decomposition for $3$-- and $4$--folds fibred by varieties with trivial Chow group of zero--cycles, J. Alg. Geom. 24 (2015), 51---80,

%\bibitem{K2} A. Vistoli, Alexander duality in intersection theory, Comp. Math. 70 (1989), 199---225,

\bibitem{Voe} V. Voevodsky, A nilpotence theorem for cycles algebraically equivalent to zero, Internat. Math. Research Notices 4 (1995), 187---198,

%\bibitem{V9} C. Voisin, Remarks on zero--cycles of self--products of varieties, in: Moduli of vector bundles, Proceedings of the Taniguchi Congress  (M. Maruyama,  editor), Marcel Dekker New York Basel Hong Kong 1994,

%\bibitem{V10} C. Voisin, Remarks on filtrations on Chow groups and the Bloch conjecture,
%Annali di matematica pura ed applicata 183 (2004), 421---438,

%\bibitem{V20} C. Voisin, Miroirs et involutions sur les surfaces K3, Ast\'erisque 218 (1993), 273---323,

%\bibitem{V12} C. Voisin, Sur les z\'ero--cycles de certaines hypersurfaces munies d'un automorphisme, Annali della Scuola Norm. Sup. di Pisa Vol. 29 (1993), 473---492,

%\bibitem{V13} C. Voisin, Chow rings and decomposition theorems for $K3$ surfaces and Calabi---Yau hypersurfaces, Geom. Topol.
%16 (2012), 433---473,

%\bibitem{V11} C. Voisin, Symplectic involutions of $K3$ surfaces act trivially on $CH_0$, Documenta Math. 17 (2012), 851---860,

%\bibitem{V0} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, Ann. Sci. Ecole Norm. Sup. 46, fascicule 3 (2013), 449---475,

%\bibitem{V1} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, II, J. Math. Sci. Univ. Tokyo  22 (2015), 491---517,

%\bibitem{V15} C. Voisin, On the universal $CH_0$ group of cubic hypersurfaces, to appear in Journal Eur. Math. Soc.,

\bibitem{V8} C. Voisin, Bloch's conjecture for Catanese and Barlow surfaces, J. Differential Geometry 97 (2014), 149---175,

\bibitem{Vo} C. Voisin, Chow Rings, Decomposition of the Diagonal, and the Topology of Families, Princeton University Press, Princeton and Oxford, 2014,

\bibitem{Xu} Z. Xu, Algebraic cycles on a generalized Kummer variety, arXiv:1506.04297v1.

\end{thebibliography}

\RequirePackage{fix-cm}
\documentclass[smallextended,envcountsame]{svjour3} 
\smartqed  
\usepackage{graphicx, amssymb}
\usepackage{times,a4wide,mathrsfs}
\usepackage{amsmath}
\usepackage{amsfonts}

\newcommand{\C}{\mathbb{C}}

\newcommand{\QQ}{\mathbb{Q}}
\newcommand{\NN}{\mathbb{N}}
\newcommand{\PP}{\mathbb{P}}
\newcommand{\LLL}{\mathbb{L}}

\newcommand{\MM}{\mathcal M}

\newcommand{\rom}{\romannumeral}

\makeatletter
\newcommand*{\da@rightarrow}{\mathchar"0\hexnumber@\symAMSa 4B }
\newcommand*{\da@leftarrow}{\mathchar"0\hexnumber@\symAMSa 4C }
\newcommand*{\xdashrightarrow}[2][]{%
  \mathrel{%
    \mathpalette{\da@xarrow{#1}{#2}{}\da@rightarrow{\,}{}}{}%
  }%
}
\newcommand{\xdashleftarrow}[2][]{%
  \mathrel{%
    \mathpalette{\da@xarrow{#1}{#2}\da@leftarrow{}{}{\,}}{}%
  }%
}
\newcommand*{\da@xarrow}[7]{%
  \sbox0{$\ifx#7\scriptstyle\scriptscriptstyle\else\scriptstyle\fi#5#1#6\m@th$}%
  \sbox2{$\ifx#7\scriptstyle\scriptscriptstyle\else\scriptstyle\fi#5#2#6\m@th$}%
  \sbox4{$#7\dabar@\m@th$}%
  \dimen@=\wd0 %
  \ifdim\wd2 >\dimen@
    \dimen@=\wd2 %
  \fi
  \count@=2 %
  \def\da@bars{\dabar@\dabar@}%
  \@whiledim\count@\wd4<\dimen@\do{%
    \advance\count@\@ne
    \expandafter\def\expandafter\da@bars\expandafter{%
      \da@bars
      \dabar@ 
    }%
  }%
  \mathrel{#3}%
  \mathrel{%
    \mathop{\da@bars}\limits
    \ifx\\#1\\%
    \else
      _{\copy0}%
    \fi
    \ifx\\#2\\%
    \else
      ^{\copy2}%
    \fi
  }%
  \mathrel{#4}%
}
\makeatother

\newcommand\undermat[2]{
  \makebox[0pt][l]{$\smash{\underbrace{\phantom{
    \begin{matrix}#2\end{matrix}}}_{\text{$#1$}}}$}#2}

\newtheorem{convention}{Conventions}

\newtheorem{nonumbering}{Theorem}

\newtheorem{nonumberingp}{Proposition}

 \journalname{Annali dell'Universita di Ferrara}

\begin{document}

\title{Some elementary examples of quartics with finite--dimensional motive}

\author{Robert Laterveer}

\institute{CNRS - IRMA, Universit\'e de Strasbourg \at
              7 rue Ren\'e Descartes \\
              67084 Strasbourg cedex\\
              France\\
              \email{laterv@math.unistra.fr}   }

\date{Received: date / Accepted: date}

\maketitle

\begin{abstract}
This small note contains some easy examples of quartic hypersurfaces that have finite--dimensional motive. As an illustration, we verify a conjecture of Voevodsky (concerning smash--equivalence) for some of these special quartics.
 \end{abstract}

\keywords{Algebraic cycles \and Chow groups \and motives \and finite--dimensional motives \and quartics} 

\subclass{14C15, 14C25, 14C30.}

\section{Introduction}

The notion of finite--dimensional motive, developed independently by Kimura and O'Sullivan \cite{Kim}, \cite{An}, \cite{MNP}, \cite{J4}, \cite{Iv} has given important new impetus to the study of algebraic cycles. To give but one example: thanks to this notion, we now know the Bloch conjecture is true for surfaces of geometric genus zero that are rationally dominated by a product of curves \cite{Kim}. It thus seems worthwhile to find concrete examples of varieties that have finite--dimensional motive, this being (at present) one of the sole means of arriving at a satisfactory understanding of Chow groups. 

The present note aims to contribute something to the list of examples of varieties with finite--dimensional motive, by considering quartic hypersurfaces.
In any dimension, there is one famous quartic known to have finite--dimensional motive: the Fermat quartic
  \[ (x_0)^4+(x_1)^4+\cdots+(x_{n+1})^4=0\ .\]
  The Fermat quartic has finite--dimensional motive because it is rationally dominated by a product of curves, and the indeterminacy locus is again of Fermat type \cite{Sh}. 
  The main result of this note presents, in any dimension, a family of quartics (containing the Fermat quartic) with finite--dimensional motive:
  
  \begin{nonumbering}[=theorem \ref{main}] The following quartics have finite--dimensional motive:

\noindent
(\rom1) a smooth quartic $X\subset\PP^{3k-1}(\C)$ given by an equation
  \[  f_1(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1})  =0\ ,\]
  where the $f_i$ define smooth quartic curves;
  
  \noindent
  (\rom2) a smooth quartic $X\subset\PP^{3k}(\C)$ given by an equation
   \[ f_1(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1}) +(x_{3k})^4 =0\ ,\]
  where the $f_i$ define smooth quartic curves;
   
  \noindent
  (\rom3) a smooth quartic $X\subset\PP^{3k+1}(\C)$ given by an equation 
   \[ f_1(x_0,x_1,x_2)+ f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1}) +h(x_{3k},x_{3k+1}) =0\ ,\]
  where the $f_i$ define smooth quartic curves;   
   \end{nonumbering}  

The proof is an elementary argument, based on the fact that the inductive structure exhibited by Shioda \cite{Sh}, \cite{KS} still applies to this kind of hypersurfaces.

To illustrate how nicely finite--dimensionality allows to understand algebraic cycles, we provide an application to a conjecture of Voevodsky's \cite{Voe}:

\begin{nonumberingp}[=proposition \ref{smash}] Let $X$ be a quartic as in theorem \ref{main}, and suppose $X$ has odd dimension. Then numerical equivalence and smash--equivalence coincide for all algebraic cycles on $X$.
\end{nonumberingp}

(For the definition of smash--equivalence, cf. definition \ref{sm}.)




\vskip0.4cm

\begin{convention} All varieties will be projective irreducible varieties over $\C$.

For smooth $X$ of dimension $n$, we will denote by $A^j(X)=A_{n-j}(X)$ the Chow group $CH^j(X)\otimes{\QQ}$ of codimension $j$ cycles under rational equivalence. The notations $A^j_{hom}(X)$ and $A^j_{AJ}(X)$ will denote the subgroup of homologically trivial resp. Abel--Jacobi trivial cycles. The category $\MM_{\rm rat}$ will denote the (contravariant) category of Chow motives with $\QQ$--coefficients over $\C$.
For a smooth projective variety, $h(X)=(X,\Delta_X,0)$ will denote its motive in $\MM_{\rm rat}$. 

The group $H^\ast(X)$ will denote singular cohomology with $\QQ$--coefficients.





\end{convention}

\section{Finite--dimensionality}

We refer to \cite{Kim}, \cite{An}, \cite{MNP}, \cite{Iv}, \cite{J4} for the definition of finite--dimensional motive. 
An essential property of varieties with finite--dimensional motive is embodied by the nilpotence theorem:

\begin{theorem}[Kimura \cite{Kim}]\label{nilp} Let $X$ be a smooth projective variety of dimension $n$ with finite--dimensional motive. Let $\Gamma\in A^n(X\times X)_{}$ be a correspondence which is numerically trivial. Then there is $N\in\NN$ such that
     \[ \Gamma^{\circ N}=0\ \ \ \ \in A^n(X\times X)_{}\ .\]
\end{theorem}

 Actually, the nilpotence property (for all powers of $X$) could serve as an alternative definition of finite--dimensional motive, as shown by a result of Jannsen \cite[Corollary 3.9]{J4}.
 
 \begin{conjecture}[Kimura \cite{Kim}]\label{findim}
  All smooth projective varieties have finite--dimensional motive.
  \end{conjecture}
  
   We are still far from knowing this, but at least there are quite a few non--trivial examples:
 
\begin{remark} 
The following varieties have finite--dimensional motive: abelian varieties, varieties dominated by products of curves \cite{Kim}, $K3$ surfaces with Picard number $19$ or $20$ \cite{P}, surfaces not of general type with $p_g=0$ \cite[Theorem 2.11]{GP}, certain surfaces of general type with $p_g=0$ \cite{GP}, \cite{PW}, \cite{V8}, Hilbert schemes of surfaces known to have finite--dimensional motive \cite{CM}, generalized Kummer varieties \cite[Remark 2.9(\rom2)]{Xu},
 threefolds with nef tangent bundle \cite{Iy} (an alternative proof is given in \cite[Example 3.16]{V3}), fourfolds with nef tangent bundle \cite{Iy2}, log--homogeneous varieties in the sense of \cite{Br} (this follows from \cite[Theorem 4.4]{Iy2}), certain threefolds of general type \cite[Section 8]{V5}, varieties of dimension $\le 3$ rationally dominated by products of curves \cite[Example 3.15]{V3}, varieties $X$ with $A^i_{AJ}(X)_{}=0$ for all $i$ \cite[Theorem 4]{V2}, products of varieties with finite--dimensional motive \cite{Kim}.
\end{remark}

\begin{remark}
It is an embarassing fact that up till now, all examples of finite-dimensional motives happen to lie in the tensor subcategory generated by Chow motives of curves, i.e. they are ``motives of abelian type'' in the sense of \cite{V3}. On the other hand, there exist many motives that lie outside this subcategory, e.g. the motive of a very general quintic hypersurface in $\PP^3$ \cite[7.6]{Del}.
\end{remark}

\section{Main}

\begin{theorem}\label{main} The following quartics have finite--dimensional motive:

\noindent
(\rom1) a smooth quartic $X\subset\PP^{3k-1}(\C)$ given by an equation
  \[  f_1(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1})  =0\ ,\]
  where the $f_i$ define smooth quartic curves;
  
  \noindent
  (\rom2) a smooth quartic $X\subset\PP^{3k}(\C)$ given by an equation
   \[ f_1(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1}) +(x_{3k})^4 =0\ ,\]
  where the $f_i$ define smooth quartic curves;
   
  \noindent
  (\rom3) a smooth quartic $X\subset\PP^{3k+1}(\C)$ given by an equation 
   \[ f_1(x_0,x_1,x_2)+ f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1}) +h(x_{3k},x_{3k+1}) =0\ ,\]
  where the $f_i$ define smooth quartic curves;   
      \end{theorem}

  \begin{proof} The proof is by induction, using Shioda's trick in the guise of the following proposition (this is \cite[Remark 1.10]{KS}):
  
  \begin{proposition}[Katsura--Shioda \cite{KS}]\label{sh} Let $Z\subset \PP^{m_1+m_2}$ be a smooth hypersurface of degree $d$ defined by an equation
    \[ g_1(x_0,\ldots,x_{m_1})+g_2(x_{m_1+1},\ldots,x_{m_1+m_2})=0\ .\]
   Let $Z_1$ resp. $Z_2$ be the smooth hypersurfaces of dimension $m_1$ resp. $m_2-1$, defined as
     \[ g_1(x_0,\ldots,x_{m_1})+y^d=0\ ,\]
     resp.
     \[ g_2(x_{m_1+1},\ldots,x_{m_1+m_2})+z^d=0\ .\]   
     Then there exists a dominant rational map
     \[ \phi\colon\ \ Z_1\times Z_2\ \dashrightarrow\ Z\ ,\]
     and the indeterminacy of $\phi$ is resolved by blowing up the locus
     \[  \Bigl( Z_1\cap (y=0)\Bigr)\times \Bigl( Z_2\cap (z=0)\Bigr)\ \subset\ Z_1\times Z_2\ .\]
     \end{proposition}
      
     The induction base is $k=1$. In this case, the quartic defined in (\rom1) has finite--dimensional motive because it is a curve; the quartic surface defined in (\rom2) has finite--dimensional motive because it is dominated by a product of curves \cite[Example 11.3]{vG}; case (\rom3) is OK because any quartic threefold $X$ has
     \[ A^\ast_{AJ}(X)=0 \]
     and as such has finite--dimensional motive \cite[Theorem 4]{V2}.
     
     Next, we suppose the theorem is true for $k-1$. 
     
     Let $X\subset\PP^{3k-1}$ be a quartic as in (\rom1). According to proposition \ref{sh}, $X$ is rationally dominated by the product
     $X_1\times S$, where $S$ is a $K3$ surface as in \cite[Example 11.3]{vG}, and $X_1$ is of type
      \[ f_1(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_{k-1}(x_{3k-6},x_{3k-5},x_{3k-4}) +y^4 =0\ .\] 
      We have seen that $S$ has finite--dimensional motive, and $X_1$ has finite--dimensional motive by induction. $X$ is dominated by the blow--up of $X_1\times S$ with center
      $Z_1\times C$, where $C\subset S$ is a curve, and $Z_1\subset X_1$ is of type
      \[       f_1(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_{k-1}(x_{3k-6},x_{3k-5},x_{3k-4}) =0\ .\]
      The product $Z_1\times C$ has finite--dimensional motive (by induction), hence so has the blow--up and so has $X$.
      
      Let $X\subset\PP^{3k}$ be a quartic as in (\rom2). Now $X$ is rationally dominated by $X_1\times X_2$, where $X_1$ is of type
      \[   f(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_{k-1}(x_{3k-6},x_{3k-5},x_{3k-4}) +y^4 =0\ ,\]    
      and $X_2$ is a fourfold of type
      \[   f_k(x_{3k-3},x_{3k-2},x_{3k-1}) +(x_{3k})^4    +z^4=0\ .\]
      Both have finite--dimensional motive by induction. To resolve the indeterminacy, we need to blow up $X_1\times X_2$ in the center of type $Z_1\times Z_2$, where
      $Z_1\subset X_1$ is of type
        \[    f(x_0,x_1,x_2)+f_2(x_3,x_4,x_5)+\cdots + f_{k-1}(x_{3k-6},x_{3k-5},x_{3k-4}) =0\ .\]    
       This $Z_1$ has finite--dimensional motive (by induction), and so does $Z_2$ (any smooth quartic threefold has finite--dimensional motive). It follows that $X$ has finite--dimensional motive.
       
       Lastly, let $X\subset\PP^{3k+1}$ be a quartic as in (\rom3). Then $X$ is rationally dominated by $X_1\times C$, where $X_1$ is of type
       \[  f(x_0,x_1,x_2)+ f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1}) +y^4 =0\ ,\]   
       and $C$ is a curve. By induction, we may suppose $X_1$ has finite--dimensional motive (because we have already treated case (\rom2)). The indeterminacy locus is of the form $Z_1\times\{\hbox{points}\}$, where $Z_1$ is of type
       \[ f(x_0,x_1,x_2)+ f_2(x_3,x_4,x_5)+\cdots + f_k(x_{3k-3},x_{3k-2},x_{3k-1}) =0\ .\]
       This $Z_1$ also has finite--dimensional motive (because we have already treated case (\rom1)), hence so does the blow--up and so does $X$.
       \end{proof}

\section{Application: Voevodsky's conjecture}

\begin{definition}[Voevodsky \cite{Voe}]\label{sm} Let $X$ be a smooth projective variety. A cycle $a\in A^r(X)$ is called {\em smash--nilpotent\/} 
if there exists $m\in\NN$ such that
  \[ \begin{array}[c]{ccc}  a^m:= &\undermat{(m\hbox{ times})}{a\times\cdots\times a}&=0\ \ \hbox{in}\  A^{mr}(X\times\cdots\times X)_{}\ .
  \end{array}\]
  \vskip0.6cm

Two cycles $a,a^\prime$ are called {\em smash--equivalent\/} if their difference $a-a^\prime$ is smash--nilpotent. We will write $A^r_\otimes(X)\subset A^r(X)$ for the subgroup of smash--nilpotent cycles.
\end{definition}

\begin{conjecture}[Voevodsky \cite{Voe}]\label{voe} Let $X$ be a smooth projective variety. Then
  \[  A^r_{num}(X)\ \subset\ A^r_\otimes(X)\ \ \ \hbox{for\ all\ }r\ .\]
  \end{conjecture}

\begin{remark} It is known \cite[Th\'eor\`eme 3.33]{An} that conjecture \ref{voe} implies (and is strictly stronger than) conjecture \ref{findim}.
\end{remark}

\begin{proposition}\label{smash} Let $X$ be a smooth quartic of the type as in theorem \ref{main}. Suppose the dimension $n$ of $X$ is odd. Then
  \[  A^r_{num}(X)\ \subset\ A^r_\otimes(X)\ \ \ \hbox{for\ all\ }r\ .\]
  \end{proposition}
  
  \begin{proof} As $X$ is a hypersurface, the K\"unneth components are algebraic and the Chow motive of $X$ decomposes
    \[ h(X)=h_n(X)\oplus \bigoplus_j \LLL(n_j)\ \ \ \hbox{in}\ \MM_{\rm rat}\ .\]
    Since $A^r_{num}\bigl(\LLL(n_j)\bigr)=0$, we have
    \[ A^r_{num}(X)= A^r_{num}\bigl(h_n(X)\bigr)\ .\]
    The motive $h_n(X)$ is oddly finite--dimensional. (Indeed, as $n$ is odd we have that the motive $\hbox{Sym}^m h_n(X)\in \MM_{\rm hom}$ is $0$ for some $m>>0$. By finite--dimensionality, the same then holds in $\MM_{\rm rat}$.)
    The proposition now follows from the following result (which is \cite[Proposition 6.1]{Kim}, and which is also applied in \cite{KSeb} where I learned this):
    
    \begin{proposition}[Kimura \cite{Kim}] Let $M\in\MM_{\rm rat}$ be oddly finite--dimensional. Then
       \[ A^r_{}(M)\ \subset\ A^r_\otimes(M)\ \ \ \hbox{for\ all\ }r\ .\]
     \end{proposition}
     
     \end{proof}

\vskip0.6cm

\begin{acknowledgements} This note is a belated echo of the Strasbourg 2014---2015 groupe de travail based on the monograph \cite{Vo}. Thanks to all the participants for the pleasant and stimulating atmosphere. 
Many thanks to Yasuyo, Kai and Len for lots of enjoyable post--work ap\'eritifs. 
\end{acknowledgements}


\end{document}